# Omni-MATH: A Universal Olympiad Level Mathematic Benchmark for Large Language Models

### v2

@@ 1 Introduction
To bridge this gap, we introduce Omni-MATH, a universal Olympiad-level benchmark specifically designed for mathematical reasoning. Our dataset exclusively focuses on text-only mathematics and encompasses an extensive collection of 4,428 competition-level problems. We curate our benchmark based on the actual selection processes of mathematical Olympiads, ensuring that our dataset includes problems from both introductory competitions and professional international contests. To further explore LLMs’ performance in multiple mathematical areas, we have established a hierarchical classification of mathematical domains, including 33 sub-domains and more than 10 distinct difficulty levels (see Figure 2), allowing for a nuanced analysis of model performance across various mathematical disciplines and levels of complexity. For better evaluation, we provide a GPT-4o-based model evaluation and Omni-Judge, an open-source verifier. Our Omni-Judge provides achieves over 91% consistency with GPT-4o and 86% consistency with human judgments, providing reliable feedback. What’s more, the open-source verifier provides a simple and effective way to evaluate the Olympiad-level mathematical problems.
We conduct a comprehensive evaluation of the currently strongest models on Omni-MATH. As shown in 3, it presents a significant challenge to the existing models. Even the most capable models OpenAI o1-mini and o1-preview, renowned for their reasoning ability, only achieve an accuracy of 60.54% and 52.55%, while the highest score among open-source models is just 36.2%. In addition, the detailed categorization of mathematical disciplines and comprehensive evaluation provide new insights into current LLMs. For instance, LLMs show a marginally greater aptitude for solving algebra, while struggling significantly with discrete mathematics. The commonly used approach for test-time scaling, Best-of-N, has proven ineffective for Olympiad-level mathematics problems. This underscores the urgent need for further research into test-time scaling techniques.
To summarize, our contribution includes:
@@
2) We introduce Omni-Judge, the first mathematical verifier designed for highly challenging problems. Despite its compact size of only 7B, Omni-Judge achieves over 90% consistency with GPT-4o, providing an efficient solution for mathematical verification.
@@ 2.1 Data Collection and Annotation
Table 2: The number of data in the construction process. ”#Data” denotes the number of data, ”#R.Filter” denotes the number of data after rule-based filtering, and ”#H.Filter” denotes the number of data after human annotation.
@@ 2.2 Difficulty Classification
After conducting instance-level difficulty classification, we organized the difficulty distribution by competition. Our analysis revealed that the actual difficulty distribution rankings shown in Figure 4 closely align with the tiered classifications we previously surveyed in Figure 5. This consistency further validates the reliability of our domain classification.
@@ 2.4 Evaluation
We discover that conducting answer-level assessments for Olympiad-level problems is a complex task due to the diverse formats of the final answers produced by models (Figure 5). This diversity makes it challenging to evaluate model outputs using fixed rules, as seen in other benchmarks, and we accordingly employ model-based evaluation for the assessment of Omni-MATH. Specifically, we leverage GPT-4o to verify whether the content generated by the tested model aligns with the standard answer. In our prompt, we provide problem, reference answer, model-generated solution, querying GPT-4o to determine whether the model solution is consistent with the reference answer. Detailed prompts and model outputs can be found in the Appendix H. Additionally, in Section 4.2, we assessed the evaluative capability of GPT-4o to ensure the accuracy of our results.
To promote public accessibility for Omni-MATH, we additionally developed an open-source evaluation model, named Omni-Judge, to assess the consistency between model solutions and reference answers at a low cost. Specifically, we first constructed a dataset for training ($\sim$17618), validation ($\sim$2200), and test ($\sim$2200) based on partial results from GPT-4o evaluation, which have no overlaps of questions with each other. Then we treat this task as a similar flow of instruction following, where the extraction of the model answer, judgment, and detailed justification are all required to be generated by Omni-Judge. We accordingly test multiple instruct models while appending a case in the context to boost the performances. It is finally evaluated on the consistency between its judges and those from GPT-4o, which surprisingly exceeds 90%, demonstrating the strong feasibility of Omni-Judge. Detailed analysis can be found in Sec 4.2 and Appendix I.
@@ 3.2 Experimental Result
The efficient test-time scaling techniques warrant further development. We have employed the common approach in the current mathematical reasoning studies, known as Best-of-N (Wang et al., 2024; Gao et al., 2024; Yang et al., 2024b), utilizing the Qwen2.5-Math-RM-72B. However, we found that scaling the inference time did not yield consistent improvements in performance. We propose two potential reasons: (1) The reward model demonstrates a limited ability to supervise tasks related to Olympic-level mathematics. The reward model can successfully improve the Qwen2.5-MATH-7b-Instruct from a total accuracy of 33.22 to 35.79, but fails with Qwen2.5-MATH-72b-Instruct. The reason might be the limitations of the reward model’s discriminative ability, which is comparable to that of Qwen2.5-MATH-72b-Instruct, making it challenging to further improve performance. (2) The policy model struggles to search the correct solutions. Moreover, with the limitations on inference tokens for OpenAI o1-preview and OpenAI o1-mini set to a maximum of 4096, which is considerably lower than the RM@256 costs, we still observe performance that greatly surpasses that of vanilla models. Thus, the development of more efficient test-time enhancement approaches remains an open area for exploration.
@@ 4.2 Reliability of Judgment Models
There are several studies utilizing model-based evaluation like GPT-4 to evaluate the correctness of answers (Fang et al., 2024; Huang et al., 2024), but whether such judgment models can consistently provide reliable feedback on Olympic-level problems remains inconclusive. In this part, we conduct a meta-evaluation experiment to assess the reliability of model-generated judgments. In detail, we first build a subset containing 100 samples and engage several graduate and doctoral students to annotate each sample as golden judgments, i.e. whether each solution is aligned with the standard answer for the given problem. Next, the predictions from judgment models are compared with the gold-standard annotations. We were surprised to find that GPT-4o acquires a 98% accuracy with human annotations, demonstrating the strong reliability of both the evaluation tool and the results in Omni-MATH. Additionally, Omni-Judge also shows considerable effectiveness with an accuracy of 86%, making it a cost-effective tool for quickly evaluating and iterating models.
Since GPT-4o shows great consistency with human annotations, we treat it as a reliable proxy of human judgment and further analyze the performance of Omni-Judge in Appendix I.
@@ H.2 Case Study of Evaluation Issue
Appendix I Analysis of Omni-Judge

### v3

@@ Abstract
[Uncaptioned image] Github Repo [GitHub Page]
[Uncaptioned image] Rule-based Repo [GitHub Page]
[Uncaptioned image] Project Page [Project Page & Leaderboard]
[Uncaptioned image] Dataset [Huggingface Dataset]
[Uncaptioned image] OmniJudge [Huggingface Model]
@@ 1 Introduction
To bridge this gap, we introduce Omni-MATH, a universal Olympiad-level benchmark specifically designed for mathematical reasoning. Our dataset exclusively focuses on text-only mathematics and encompasses an extensive collection of 4,428 competition-level problems. We curate our benchmark based on the actual selection processes of mathematical Olympiads, ensuring that our dataset includes problems from both introductory competitions and professional international contests. To further explore LLMs’ performance in multiple mathematical areas, we have established a hierarchical classification of mathematical domains, including 33 sub-domains and more than 10 distinct difficulty levels (see Figure 2), allowing for a nuanced analysis of model performance across various mathematical disciplines and levels of complexity. For better evaluation, we provide a GPT-4o-based model evaluation and an open-source verifier, Omni-Judge, besides the traditional rule-based evaluation. Our Omni-Judge provides achieves over 91% consistency with GPT-4o and 86% consistency with human judgments, providing reliable feedback. What’s more, the open-source verifier provides a simple and effective way to evaluate the Olympiad-level mathematical problems.
We conduct a comprehensive evaluation of the currently strongest models on Omni-MATH. As shown in Table 3, it presents a significant challenge to the existing models. Even the most capable models OpenAI o1-mini and o1-preview, renowned for their reasoning ability, only achieve an accuracy of 60.54% and 52.55%, while the highest score among open-source models is just 36.2%. In addition, the detailed categorization of mathematical disciplines and comprehensive evaluation provide new insights into current LLMs. For instance, LLMs show a marginally greater aptitude for solving algebra, while struggling significantly with discrete mathematics. The commonly used approach for test-time scaling, Best-of-N, has proven ineffective for Olympiad-level mathematics problems. This underscores the urgent need for further research into test-time scaling techniques. To summarize, our contribution includes:
@@
2) We comprehensively examine different evaluation models. Our experiments reveal that GPT-4o evaluation can align well with human evaluation with an accuracy of 98% and the Omni-Judge achieves over 90% consistency with GPT-4o, providing an efficient and reliable evaluation.
@@ 2.1 Data Collection and Annotation
Data Annotation
Table 2: The number of data in the construction process. ”#Data” denotes the number of data, ”#R.Filter” denotes the number of data after rule-based filtering, and ”#H.Filter” denotes the number of data after human-annotation.
@@ 2.2 Difficulty Classification
After conducting instance-level difficulty classification, we organized the difficulty distribution by competition. Our analysis revealed that the actual difficulty distribution rankings shown in Figure 4 closely align with the tiered classifications we previously surveyed in Table 5. This consistency further validates the reliability of our domain classification.
@@ 2.4 Evaluation
We discover that conducting answer-level assessments for Olympiad-level problems is a complex task due to the diverse formats of the final answers produced by models (Figure 5 and Section H.3). This diversity makes it challenging to evaluate model outputs using fixed rules, as seen in other benchmarks, and we accordingly employ model-based evaluation for the assessment of Omni-MATH. Specifically, we leverage GPT-4o to verify whether the content generated by the tested model aligns with the standard answer. In our prompt, we provide problem, reference answer, model-generated solution, querying GPT-4o to determine whether the model solution is consistent with the reference answer. Detailed prompts and model outputs can be found in the Appendix H. Additionally, in Section 4.2, we assessed the evaluation capability of GPT-4o to ensure the accuracy of our results.
To promote public accessibility for Omni-MATH, we additionally filtered a subset named Omni-MATH-Rule for reliable rule-based evaluation, as in Appendix K. What’s more, for total set evaluation, we also developed an open-source evaluation model, named Omni-Judge, to assess the consistency between model solutions and reference answers at a low cost. Specifically, we first constructed a dataset for training ($\sim$17618), validation ($\sim$2200), and test ($\sim$2200) based on evaluation results from GPT-4o, which have no overlaps of questions with each other. Then we treat this task as a similar flow of instruction following, where the extraction of the model answer, judgment, and detailed justification are all required to be generated by Omni-Judge. We accordingly test multiple instruct models while appending a case in the context to boost the performances. It is finally evaluated on the consistency between its judges and those from GPT-4o, which surprisingly exceeds 90%, demonstrating the strong feasibility of Omni-Judge. Detailed analysis can be found in Sec 4.2 and Appendix J.
@@ 3.2 Experimental Result
The efficient test-time scaling techniques warrant further development. We have employed the common approach in the current mathematical reasoning studies, known as Best-of-N (Wang et al., 2024; Gao et al., 2024; Yang et al., 2024b), utilizing the Qwen2.5-Math-RM-72B. However, we found that scaling the inference time did not yield consistent improvements in performance. We propose two potential reasons: (1) The reward model demonstrates a limited ability to supervise tasks related to Olympic-level mathematics. (2) The policy model struggles to search for the correct solutions. Moreover, with the limitations on inference tokens for OpenAI o1-preview and OpenAI o1-mini set to a maximum of 4096, which is considerably lower than the RM@256 costs, we still observe performance that greatly surpasses that of vanilla models. Thus, the development of more efficient test-time enhancement approaches remains an open area for exploration. The further discussion about the underlying problem is detailed in Appendix I.
@@ 4.2 Reliability of Judgment Models
Despite the traditional rule-based evaluation, model-based evaluation provides a flexible evaluation in an end-to-end manner without the need to extract the answers and build complex rules. There are several studies utilizing model-based evaluation like GPT-4 to evaluate the correctness of answers (Fang et al., 2024; Huang et al., 2024), but whether such judgment models can consistently provide reliable feedback on Olympic-level problems remains inconclusive. In this part, we conduct a meta-evaluation experiment to assess the reliability of model-generated judgments. In detail, we first build a subset containing 100 samples and engage several graduate and doctoral students to annotate each sample as golden judgments, i.e. whether each solution is aligned with the standard answer for the given problem, as details in Appendix C.3. Next, the predictions from judgment models are compared with the gold-standard annotations. We were surprised to find that GPT-4o acquires a 98% accuracy with human annotations, demonstrating the strong reliability of both the evaluation tool and the results in Omni-MATH. Additionally, Omni-Judge also shows considerable effectiveness with an accuracy of 86%, making it a cost-effective tool for quickly evaluating and iterating models.
Since GPT-4o shows great consistency with human annotations, we treat it as a reliable proxy of human judgment and further analyze the performance of Omni-Judge in Appendix J.
@@ C.2 Annotation Details
C.3 Details of Meta-Evaluation Set
The 100 samples of the meta-evaluation set mentioned in Section 4.2 were derived from the inference results of DeepSeek-Coder-V2, GPT-4o, and Qwen2.5-MATH-72b-Instruct, sampled according to difficulty to ensure representation across all levels of difficulty without overlapping questions. Although the inference results come from different models, our ultimate evaluation target is the human alignment of judgment models—GPT-4o and Omni-Judge. Therefore, all 100 examples are designed for GPT-4o or Omni-Judge to provide reliable human evaluation.
Our experiments revealed that GPT-4o achieved highly accurate results under conditions of high agreement among human annotators (97% inter-agreement) given the solution generated by 3 different models. For annotation, we employed two PhD students and two Master’s students, with each annotator labeling 50 problems. Moreover, they were instructed to determine whether the model’s answers were correct based mostly on the correct answers rather than their own judgments. Each annotator was responsible for 50 questions, allowing us to ensure that each question was annotated by two annotators, facilitating cross-validation to assess inter-agreement. During individual annotation, three out of 100 instances of inconsistent results were observed; however, after discussion, the annotators reached a consensus on the answers. The discrepancies primarily stemmed from GPT-4o misinterpreting the ”Student Answer” as the ”Ground Truth” under the current prompt, which we attribute to prompt formatting issues. This problem was resolved by adjusting the few-shot prompt format.
@@ H.2 Case Study of Evaluation Issue
Table 7: Distributions of answer formats in 200 problems randomly sampled from Omni-MATH.
Number LaTex Tuple Multi-LaTex Function Multi-Function Text
95 51 2 8 9 1 34
H.3 Distributions of Answer Types
Rule-based answer extraction and evaluation were typical methods in LLM assessment of mathematical capacity, as utilized in prior works like MATH Hendrycks et al. (2021); He et al. (2024); Huang et al. (2024). Nevertheless, real-world mathematical problems vary widely in domain and corresponding answer format. To better assess the capabilities of LLMs as real-world problem solvers, Omni-MATH goes beyond simple types of answer formats like multiple-choice or direct number, incorporating more problems with diverse formats.
In this part, we conduct an analysis from this perspective with randomly sampled 200 problems, as shown in Table 7, where the answer formats are split into several types:
  • Number A numeric answer.
  • LaTex A LaTeX formula.
  • Tuple A tuple-formatted answer.
  • Multi-LaTex A combination of multiple LaTeX formulas.
  • Function A custom-defined function.
  • Multi-Function A combination of multiple custom-defined functions.
  • Text:  A free-form explanation.
From the statistics, problems formatted in Number, LaTex and Tuple, which can easily extracted with rules, account for only 74% of the total, while problems with more complex answers also make up a significant portion, suggesting that results across these problems is crucial. To this end, the GPT-4 evaluation applied in Omni-MATH can be useful and accurate for it effectively covers different formats and achieves an impressive 98% consistency with human annotation (see Sec 4.2). It ensures that Omni-MATH offers a thorough and fair assessment of mathematical capacity.
Appendix I Further Discussion on Test-time Scaling Results
Table 3 indicates that RM@256 fails to perform better than RM@8 on Qwen2.5-MATH-72b-Instruct, which seems counterintuitive. We find that it can be more likely attributed to the fact that the RM involved has insufficient supervision over Olympiad-level reasoning tasks than policy models. Specifically, it cannot accurately select the correct reasoning path from numerous candidates. Here we demonstrate it with the following test.
We measured the performance of the policy model: Qwen2.5-MATH-Instruct across N samples. If at least one reasoning trace is correct, we consider the example to assess the interference in RM selection, we also measured the proportion of correct COTs within the passing examples. Considering the inference cost, we sampled a subset of 500 instances according to difficulty to ensure that the distribution of problem difficulty is similar to the original dataset. The experimental result is shown below:
Table 8: The performance of the policy model under Pass@K and the proportion of correct COT.
Metric Pass@1 Pass@8 Pass@16 Pass@32
Acc 40.8% 57.6% 63% 67.2%
Correct COT Proportion 100% 68.5% 62.5% 58.5%
The results show that as the inference sampling increases, the policy model is capable of solving the majority of the problems. However, even after 32 samples, 33.8% of the problems still do not yield the correct solution. The second row of the table illustrates that as the number of samples increases, the proportion of interference items also increases. In the case of 32 samples, the correct answer proportion of the passed cases drops from 100% to 58.5%. This indicates that as the sampling count increases, it becomes increasingly challenging for the reward model to select the correct COT.
@@ Appendix I Further Discussion on Test-time Scaling Results
Appendix J Analysis of Omni-Judge
@@
Appendix K The construction details of Omni-MATH-Rule
As mentioned in Section 2.4, we additionally filtered a subset named Omni-MATH-Rule for efficiently reliable rule-based evaluation.
We began by rewriting a new version of the rule-matching code using the Qwen2.5-math rule evaluation repository, primarily utilizing SymPy. Then we performed an initial filtering of the entire dataset of problems, which was done as follows:
We first applied the rule evaluation to the results from O1-mini, ensuring that all answers were within the ”boxed” format using the system prompt. We then extracted the answers using the ”last_boxed” method, followed by a rule evaluation, which was fully consistent with the MATH dataset.
We then compared the rule-based and GPT-4o-based results (given that GPT-4o-based evaluation can perform over 98% consistency with human evaluation). Based on their judgments of the policy model generation, we identified the following cases:
  • Rule-based Correct and GPT-4o-based Correct: This indicates that the answer was accepted by both Rule-based and GPT-4o. We consider this case to have a high probability of evaluation correctness, and we labeled it as a ”positive case.”
  • Rule-based False and GPT-4o-based False: This indicates that both Rule-based and GPT-4o rejected the answer. However, the rule-based rejection could be due to generalization issues, meaning it may not align with GPT-4o’s rejection. We label this as an ”uncertain case,” and we need to use strict rules to filter out only those cases that can be reliably evaluated.
  • Rule-based and GPT-4o-based judgments inconsistent: This case could suggest evaluation problems with either the GPT-4o evaluation or the rule evaluation. Therefore, we marked this as an ”inconsistent case.”
The statistics of the cases above can be shown as follows:
Table 10: statistics of Rule-based and GPT-4o-based Results
Rule-based GPT-4o-based Count
Correct Correct 2053
Correct False 66
False Correct 545
False False 1764
We found that less than 2% of cases (fewer than 66 out of 4428) were identified by GPT-4o as incorrect answers while still matching the rules correctly. We believe these instances reliably indicate evaluation errors by GPT-4o. Upon careful examination, we discovered that some cases experienced parsing failures, resulting in both the predicted and ground truth outputs being matched as empty strings, which accounted for 30 out of the 66 cases. After identifying this issue, we removed these cases from the testable dataset. Consequently, the actual number of instances categorized as ”Rule-based: Correct and GPT-4o-based: False” is reduced to 36. Among these, several remain contentious; for example, GPT-4o indicated that a condition was omitted in the answer.
We then added the ”positive cases” and ”uncertain cases” to the testable set which indicates the cases are reliable with rule-based evaluation. The ”inconsistent cases” were placed directly into the untestable set.
We applied strict rules (using regular expressions) to filter the testable set, removing cases that were more complex, such as open-ended text, multi-part LaTeX, function-based problems, or multiple functions, retaining only simpler cases that the rules could match effectively, which can filter reliable cases in ”uncertain case” as much as possible.
This filtering process resulted in 2925 problems suitable for rule-based evaluation. We then employed two PhD students to annotate the rule-based evaluated data. The task was to classify the type of answer based on the aforementioned experiments (number or tuples or latex and so on) and determine whether it could be evaluated by the rule system. We ensured that each problem was annotated by two students, and our cross-validation accuracy was 98% (2869/2925). Inconsistent cases were placed in the untestable set. This annotation process removed some tuples that were still unsuitable for evaluation and some complex LaTeX cases. Ultimately, We ended up with a testable set of 2821 problems and an untestable set of 1607 problems. We analyzed the difficulty distribution of the testable set, which is basically consistent with the distribution of our entire dataset.
To further validate the reliability of the testable set and the rule evaluation, we evaluated some models on the current subset. The rankings and the acc of the models were consistent with the evaluation results from GPT-4o.
Table 11: The leaderboard of total accuracy of Omni-MATH-Rule using rule-based evaluation and Omni-MATH using GPT-4o evaluation.
Model Acc @Rule 2821 Acc @GPT-4o 4428
o1-mini 62.2% 60.54%
o1-preview 51.7% 52.55%
qwen2.5-MATH-72b-Instruct 35.7% 36.20%
qwen2.5-MATH-7b-Instruct 32.3% 33.22%
GPT-4o 29.2% 30.49%
NuminaMATH-72b-cot 27.1% 28.45%
DeepseekMATH-7b-RL 14.9% 16.12%
Appendix L Data Examples of Omni-MATH
Figure 18: An example from China National Mathematical Olympiad (CMO) in Omni-MATH.
Figure 19: An example from USA Mathematical Olympiad (USAMO) in Omni-MATH.